OpenStudy (anonymous):

a cream is sold in 26-gram container. The average amount cream used per application is 2 3/5 grams. How many applications can be made with the container?

OpenStudy (anonymous):

\[{{Total Weight}\over{One Time Application}}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

25/ 2 3/5?

OpenStudy (anonymous):

26*

OpenStudy (ash2326):

Cream is sold in a 26 gram container, which means there is 26 gram of cream in the container, Average amount of cream used per application= \(\large 2 \frac{3}{5}=2+ \frac{3}{5}=\frac{10+3}{5}=\frac{13}{5} grams\) No. of applications= \[\frac{26}{\frac{13}{5}}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

143/5?

OpenStudy (anonymous):

how it can be 143/5

OpenStudy (anonymous):

becuz 26/1+13/5= 143/5

OpenStudy (anonymous):

Total = 26 One time = 13/5 So No. of times = \[{26}\over{{13}\over{5}}\]=\[{{26\times5}}\over{13}\]=10 times

OpenStudy (anonymous):

so its 10?

OpenStudy (anonymous):

yes